# Why We Need More than General Multimodal LLMs

### v1

@@ 6 A Call on Specialized Spatial-Temporal Designs for Music AVQA
(f) Audio-Visual Location QA.
(g) Audio-Visual Comparative QA.
(h) Audio-Visual Temporal QA.
(i) Audio-Visual QA Average.
(a) Methods on Music-AVQA 9879157 .
(b) Methods on Music-AVQA-R ma2024look .
@@
Figure 3: Radar plots showing the per-type average accuracy of model groups with and without spatial-temporal design across 13 QA categories on (a) Music-AVQA 9879157 and (b) Music-AVQA-R ma2024look . Each axis corresponds to a QA type spanning audio, visual, and audio-visual reasoning, including the overall average (Total-Average). The filled  green polygon in Figure 3(a) and  purple polygon in Figure 3(b) represent the mean accuracy across QA types for models with spatial-temporal design, while the blue polygon represents the average performance of models without such design. Models with spatial-temporal design consistently achieve higher accuracy across all modality groups. These advantages persist under distribution shift in the robustness-focused Music-AVQA-R dataset.
@@
Spatial-temporal design enhances audio QA by supporting fine-grained tracking of overlapping sources and temporally evolving acoustic cues. Audio-related questions in Music AVQA—such as instrument counting or loudness comparison—require models to distinguish simultaneous sound sources, localize temporal onsets, and resolve dynamic variations across time. As shown in Figures 2(a) and 2(b), models with spatial-temporal design consistently outperform others. LAST-Att 10484352 achieves the highest audio counting accuracy at 85.71%, benefiting from repeated cross-attention between question-guided Swin-Transformer features and spectrogram patches from an Audio Spectrogram Transformer, which helps the model focus on musically salient moments. Amuse diao2024learning , with 83.58% average audio QA accuracy, aligns audio-video streams using beat-synchronous features and temporally-adaptive fusion modules, allowing it to isolate relevant auditory content even under polyphonic conditions. DG-SCT duan2023cross further introduces bidirectional attention layers across temporal, spatial, and channel dimensions, dynamically adjusting audio-visual focus based on the question’s semantics. By contrast, models lacking spatial-temporal structure—such as MCAN (67.47%) and CONVLSTM (66.73%)—often rely on global feature pooling or frame-agnostic fusion, making them vulnerable to overlap, misalignment, and temporal drift. Notably, spatial-temporal designs adopt recurring architectural motifs: temporal segment selection (PSTP-Net li2023progressive , AVST 9879157 ), audio-guided visual attention (DG-SCT, LSTTA liu2023parameter ), and fine-grained cross-modal alignment (Meerkat 10.1007/978-3-031-73039-9_4 ). These mechanisms are well-suited for modeling music’s complex structure, where overlapping instruments and evolving rhythms require localized reasoning in both time and space. The strong performance of spatial-temporal models across audio QA tasks confirms their value in resolving multi-instrument scenarios and detecting temporally grounded acoustic attributes.
@@
Spatial-temporal design improves visual QA by enhancing spatial disambiguation and capturing motion cues over time. Visual-related questions in Music AVQA—such as counting instruments or identifying positions—often involve tracking multiple performers, detecting visual cues of articulation (e.g., bowing, striking), and resolving spatial relationships within densely packed frames. As shown in Figures 2(c)2(e), models with spatial-temporal components generally achieve stronger accuracy. For example, LSTTA liu2023parameter (82.03% visual QA average) combines short-term semantic interaction and long-term semantic filtering modules to capture both local gestures and global scene dynamics, enabling precise reasoning about when and where instruments are engaged. DG-SCT duan2023cross (82.08%) uses cross-modal temporal attention guided by audio prompts to enhance visual token selection, focusing on visually active regions corresponding to sounding instruments. PSTP-Net li2023progressive (77.26%) implements a region refinement module that explicitly filters visual patches within question-relevant segments, improving spatial disambiguation. While spatial-temporal modeling is effective, some models without it still perform competitively—most notably CAT ye2024cat (86.10%), which leverages large-scale pretrained vision encoders (ImageBind) and LLaMA2 to infer structure implicitly. However, such models may rely heavily on correlation learned from pretraining, rather than explicit reasoning about visual dynamics. Spatial-temporal models, by contrast, explicitly model the temporal unfolding of gestures and the spatial focus of performer activity—important properties in musical scenes where instrument positions are static but their activation varies over time. These architectural patterns help stabilize attention and reduce confusion when multiple instruments are visually present but only some are active, contributing to more consistent visual QA performance across counting and localization tasks.
@@
Spatial-temporal design is critical for audio-visual QA, where accurate reasoning requires precise temporal and spatial alignment between modalities. Among all Music AVQA categories, audio-visual questions impose the strongest demand on cross-modal synchronization, requiring the model to associate specific acoustic events with their visual sources over time. As shown in Figures 2(f)2(i) and Table 2, models with spatial-temporal components consistently achieve higher accuracy across AV-Existential, AV-Counting, AV-Location, AV-Comparative, and AV-Temporal types. Amuse diao2024learning reaches 82.43% on overall AV questions by leveraging segment-level alignment between synchronized beat-level audio and video inputs and applying cross-modal adapters at each step. PSTP-Net li2023progressive adopts a progressive three-stage pipeline: temporal segment selection, spatial region refinement, and audio-guided attention, resulting in 72.57% AV average. Meerkat 10.1007/978-3-031-73039-9_4 further enhances local alignment by explicitly modeling cross-modal transport between audio patches and visual regions, and enforces bounding box constraints for grounding, yielding strong performance on AV-Comparative and AV-Location. In contrast, models without spatial-temporal design—such as MCAN (57.80%), GPT-4o (50.08%), and Qwen2.5-VL (47.75%)—struggle to resolve fine-grained multimodal relationships. While CAT ye2024cat achieves 83.20% AV average through large-scale pretrained encoders, its performance drops on AV-Temporal and AV-Location tasks that require precise temporal ordering or spatial binding. These results support that spatial-temporal designs—especially those involving temporally segmented reasoning, audio-guided spatial focus, and per-frame fusion—enable the model to track which instrument is sounding, when, and in which location, which is critical for answering questions such as “Did the cello on the left play after the drum on the right?”. Without such structure, models tend to conflate co-occurring signals or miss temporally offset actions, leading to lower accuracy in complex cross-modal scenarios.
@@ 7 A Call on Specialized Musical Designs for Music AVQA
Structured musical features—such as tempo, key, downbeats, and chord progressions—can provide a coherent framework for aligning audio-visual inputs across time. These symbolic or MIR-derived signals offer interpretable, temporally smooth trajectories that reflect the hierarchical organization of music, such as phrases, sections, and transitions. Crucially, they abstract away from low-level waveform fluctuations and offer a musically meaningful scaffold that persists across different genres, tempos, and instrumentation. By integrating them as auxiliary inputs or attention-guiding signals, models may improve their ability to capture long-range dependencies, maintain rhythmic continuity, and resolve ambiguous instrument interactions—especially in polyphonic or ensemble contexts. This structured conditioning can serve as a musical inductive bias, particularly helpful in complex multimodal scenes where overlapping sources challenge simple bottom-up fusion strategies, and where salient events may not be visually or acoustically distinct without temporal alignment cues.
@@
Many Music AVQA questions require reasoning over implicit causal or temporal relationships—for example, identifying which performer initiated a musical phrase, or determining whether an instrument’s entrance shifted the ensemble’s dynamic balance. These questions often lack explicit step-level supervision, making it difficult to learn reasoning paths from labels alone. To address this, models can incorporate latent reasoning trajectories: structured internal variables that represent evolving hypotheses about the musical scene. Rather than directly mapping inputs to answers, the model infers intermediate latent states—such as “which instrument is currently leading,” “how the rhythmic intensity is changing,” or “which performer is preparing to enter”—and updates these states over time as more multimodal evidence arrives. Architecturally, this can be implemented via hierarchical latent-variable models or recurrent variational modules, where latent states encode musical intentions, transitions, or causal flow. These hidden trajectories allow the model to simulate plausible sequences of musical events, enabling it to answer questions that require extrapolating or filling in missing links between observed signals. Crucially, this style of latent reasoning supports robust generalization by embedding inductive structure aligned with how humans infer musical cause and progression—not just surface-level audio-visual co-occurrence.
@@
Some musical questions—especially those involving temporal or causal dependencies—require sequential sub-decisions to reach the correct answer. For instance, the question “Which instrument enters after the piano stops?” involves: (1) detecting piano cessation, (2) identifying subsequent onsets, and (3) selecting the earliest new instrument. Rather than treating such questions as black-box classification, models can be explicitly trained to emit intermediate reasoning steps, either through supervised rationales or pseudo-labels derived from MIR-based event detection. This approach—akin to chain-of-thought (CoT) prompting in LLMs—improves transparency, encourages modular subgoal learning, and helps the model maintain alignment across modalities. Moreover, step-wise supervision can highlight failure points in temporal or semantic inference, offering clearer diagnostics for model improvement. In music contexts, CoT chains can incorporate domain-specific steps such as beat alignment, timbre matching, or onset-event attribution. These interpretable intermediate traces not only support higher accuracy on multi-stage queries but also make it easier to identify reasoning shortcuts and dataset biases.
@@ 8 Conclusion
This paper underscores the limitations of general-purpose MLLMs for domain-specific tasks such as Music AVQA. Successful Music AVQA necessitates specialized designs tailored to musical content’s unique demands: fine-grained audio-visual processing, precise temporal modeling, and integrated domain knowledge. Our central position is that while general models advance, specialized domains require tailored solutions. We call on the research community to: 1) develop more nuanced music understanding benchmarks, and 2) explore hybrid architectures combining MLLM strengths with music-specific components. The future of effective multimodal AI lies not in a universal approach, but in the thoughtful integration of general capabilities with deep, domain-specific expertise, benefiting music understanding and other complex fields.
@@ Appendix A Quantitative Comparison on Music AVQA Datasets
We present comprehensive quantitative comparisons of recent state-of-the-art methods on multiple Music AVQA datasets [69, 80, 86], shown in Table 23, and 4. We evaluate the models across a diverse set of question categories, spanning Audio-related, Visual-related, and Audio&Visual-related reasoning tasks. For each dataset, we report accuracy metrics for subcategories such as Counting, Comparative, Location, Existential, and Temporal reasoning, along with average accuracy within each modality and the overall performance.
Table 2: Comparison with state-of-the-art methods on the Music AVQA [69] test set. We report the accuracy for Audio (Counting, Comparative), Visual (Counting, Location), and Audio-Visual (Existential, Counting, Location, Comparative, Temporal) question types, along with the average accuracy for Audio, Visual, Audio-Visual, and overall. Methods highlighted with a gray background incorporate spatial-temporal designs.
@@
Amuse [24] 84.61 82.45 83.58 87.41 84.39 85.84 86.95 85.49 73.01 82.98 83.06 82.43 83.52
Audio Flamingo [61] - - - - - - - - - - - - -
AVMoE [14] - - 77.60 - - 82.70 - - - - - 71.90 75.70
AVSD [96] 72.41 61.90 68.52 67.39 74.19 70.83 81.61 58.79 63.89 61.52 61.41 65.49 67.44
AVSiam [75] - - - - - - - - - - - - -
AVST [69] 77.78 67.17 73.87 73.52 75.27 74.40 82.49 69.88 64.24 64.67 65.82 69.53 71.59
CAT [119] - - 84.90 - - 86.10 - - - - - 83.20 84.30
ChatBridge [130] - - 28.90 - - 33.10 - - - - - 43.00 78.90
CIGN [88] - - - - - - - - - - - - -
COCA [63] 79.94 67.68 75.42 75.10 75.43 75.23 83.50 66.63 69.72 64.12 65.57 69.96 72.33
CONVLSTM [32] 68.88 63.06 66.73 64.89 58.55 61.68 82.81 55.99 61.30 53.45 54.73 61.75 62.61
CrossMAE [42] - - - - - - - - - - - - -
DCL [84] - - - - - - - - - - - - -
DG-SCT [27] 83.27 64.56 76.34 81.57 82.57 82.08 81.61 72.84 65.91 64.22 67.48 70.56 74.62
EEMC [108] - - - - - - - - - - - - -
FCNLSTM [32] 69.96 61.06 66.67 63.89 58.14 60.98 83.42 56.31 60.28 50.85 56.92 61.46 62.25
GPT-4o [53] 65.42 36.07 50.75 72.36 62.30 67.33 56.12 54.84 59.23 37.84 42.35 50.08 54.06
GRU [4] 71.82 58.90 67.04 66.06 71.82 68.97 81.41 60.30 62.32 56.23 61.89 64.26 66.00
HCRN [65] 70.21 45.62 61.14 62.41 51.51 56.90 52.94 42.07 54.70 50.59 33.33 48.41 52.54
LAST-Att [80] 85.71 63.10 - 83.86 83.09 - 76.47 76.20 68.91 65.60 66.75 - 75.45
LAVisH [76] 75.59 84.13 76.86 77.45 72.91 76.29 71.91 77.52 75.81 76.75 77.62 76.31 76.10
LAViT [123] 74.36 64.56 70.73 69.39 75.65 72.56 81.21 59.33 64.91 64.22 63.23 66.64 68.93
LSTTA [77] 81.75 82.04 81.90 81.82 82.23 82.03 83.46 79.11 78.23 78.02 79.32 79.63 81.19
MAVEN [85] 79.44 54.10 72.79 80.49 93.50 86.99 87.00 66.67 73.85 54.95 68.24 69.94 74.60
MCAN [121] 75.05 54.58 67.47 68.06 72.15 70.13 81.91 54.15 53.45 52.11 47.21 57.80 62.77
MCCD [86] 83.87 71.04 79.14 79.78 76.73 78.24 80.87 51.63 71.46 64.67 64.60 67.13 72.20
Meerkat [17] - - - - - - - 85.70 - 75.98 - - -
OneLLM [43] - - - - - - - - - - - - 47.60
OPM [110] - - - - - - - - - - - - 70.80
PSAC [71] 71.33 56.07 65.68 65.89 72.07 69.02 78.59 54.80 63.11 55.96 61.17 62.75 64.92
PSTP-Net [68] 73.97 65.59 70.90 77.15 77.36 77.26 76.18 73.23 71.80 71.19 69.00 72.57 73.52
QaP [74] - - - - - - - - - - - - -
Qwen2.5-VL [5] 48.60 55.00 51.80 55.28 53.66 54.47 44.00 52.17 63.57 37.84 41.18 47.75 50.14
RefAtomNet [89] - - - - - - - - - - - - -
VALOR [78] - - 68.70 - - 74.20 - - - - - 75.30 78.90
VAST [12] 78.18 67.05 74.06 71.56 76.38 74.00 81.81 64.51 70.80 66.01 63.23 69.54 71.52
VideoLLaMA-2 [15] 79.44 52.46 69.64 81.30 82.93 82.11 77.00 63.44 77.69 59.46 64.71 68.98 72.56
VITA [34] 59.81 45.90 54.76 50.41 34.96 42.68 54.00 49.46 46.92 27.93 41.18 43.74 45.44
@@
Table 3: Comparison with state-of-the-art methods on the Music AVQA v2.0 [80] test set. We report the accuracy for Audio (Counting, Comparative), Visual (Counting, Location), and Audio-Visual (Existential, Counting, Location, Comparative, Temporal) question types, along with the average accuracy for Audio, Visual, Audio-Visual, and overall. Methods highlighted with a gray background incorporate spatial-temporal designs.
@@
Amuse [24] 84.76 83.88 84.34 88.15 85.16 86.74 88.30 87.47 78.77 84.41 85.38 85.51 85.16
AVST [69] 81.74 62.11 72.46 79.08 77.64 78.40 72.12 69.03 65.05 63.98 60.57 66.26 71.08
DG-SCT [27] 83.66 62.47 73.64 82.05 82.97 82.48 83.43 72.70 64.65 64.78 67.34 70.38 74.08
LAST-Att [80] 86.03 62.52 - 84.12 84.01 - 76.21 75.23 68.91 65.60 60.60 - 75.44
LAVisH [76] 84.36 58.57 72.17 83.25 81.46 82.40 73.26 73.45 65.64 64.26 60.82 67.75 72.34
Table 4: Comparison with state-of-the-art methods on the Music-AVQA-R [86] test set. We report the accuracy for Audio (Counting, Comparative), Visual (Counting, Location), and Audio-Visual (Existential, Counting, Location, Comparative, Temporal) question types, along with the average accuracy for Audio, Visual, Audio-Visual, and overall. Methods highlighted with a gray background incorporate spatial-temporal designs.

### v2

@@ 6 A Call on Specialized Spatial-Temporal Designs for Music AVQA
(f) AV Location QA.
(g) AV Comparative QA.
(h) AV Temporal QA.
(i) AV QA Average.
Figure 2: Accuracy comparison of Music AVQA models across representative question types, grouped by modality: (a–b) Audio, (c–e) Visual, and (f–i) Audio-Visual. Each bar corresponds to a model and is color-coded based on whether it incorporates spatial-temporal design: bars in  green,  purple, and  orange represent models that apply spatial-temporal modeling to Audio-related, Visual-related, and Audio-Visual-related question answering, respectively; bars in  blue represent models without spatial-temporal design. Across most categories, models with spatial-temporal components tend to perform more accurately, particularly on tasks requiring temporal reasoning or spatial localization.
(a) Methods on Music-AVQA.
(b) Methods on Music-AVQA-R.
@@
Figure 3: Radar plots showing the per-type average accuracy of model groups with and without spatial-temporal design across 13 QA categories on (a) Music-AVQA Li et al. (2022) and (b) Music-AVQA-R Ma et al. (2024). Each axis corresponds to a QA type spanning audio, visual, and audio-visual reasoning, including the overall average (Total-Average). The filled  green polygon in Figure 3(a) and  purple polygon in Figure 3(b) represent the mean accuracy across QA types for models with spatial-temporal design, while the blue polygon represents the average performance of models without such design. Models with spatial-temporal design consistently achieve higher accuracy.
@@
Spatial-temporal design enhances audio QA by supporting fine-grained tracking of overlapping sources and temporal acoustic variation. Audio-related questions in Music AVQA—such as instrument counting or loudness comparison—require models to distinguish simultaneous sound sources, localize temporal onsets, and resolve temporal variation. As shown in Figures 2(a) and 2(b), models with spatial-temporal design consistently outperform others. LAST-Att achieves the highest audio counting accuracy at 85.71%, using repeated cross-attention between question-guided Swin-Transformer features and spectrogram patches from an Audio Spectrogram Transformer to focus on musically salient moments. Amuse, with 83.58% average audio QA accuracy, aligns audio-video streams using beat-synchronous features and temporally-adaptive fusion modules, allowing it to isolate relevant auditory content even under polyphonic conditions. DG-SCT further introduces bidirectional attention layers across temporal, spatial, and channel dimensions, dynamically adjusting audio-visual focus by question semantics. By contrast, models lacking spatial-temporal structure—such as MCAN (67.47%) and CONVLSTM (66.73%)—often rely on global feature pooling or frame-agnostic fusion, making them vulnerable to overlap, misalignment, and temporal drift. Notably, spatial-temporal designs adopt recurring architectural motifs: temporal segment selection (PSTP-Net, AVST), audio-guided visual attention (DG-SCT, LSTTA), and fine-grained cross-modal alignment (Meerkat). These mechanisms suit for modeling music’s complex structure, where overlapping instruments and evolving rhythms require localized reasoning in both time and space. The strong performance of spatial-temporal models across audio QA tasks confirms their value in resolving multi-instrument scenarios and detecting temporally grounded acoustic attributes.
@@
Spatial-temporal design improves visual QA by enhancing spatial disambiguation and capturing motion cues over time. Visual-related questions in Music AVQA—such as counting instruments or identifying positions—often involve tracking multiple performers, detecting visual cues of articulation (e.g., bowing, striking), and resolving spatial relationships within densely packed frames. As shown in Figures 2(c)2(e), models with spatial-temporal components generally achieve stronger accuracy. For example, LSTTA (82.03% visual QA average) combines short-term semantic interaction and long-term semantic filtering modules to capture both local gestures and global scene dynamics, enabling precise reasoning about when and where instruments are engaged. DG-SCT (82.08%) uses cross-modal temporal attention guided by audio prompts to enhance visual token selection, focusing on visually active regions corresponding to sounding instruments. PSTP-Net (77.26%) implements a region refinement module that explicitly filters visual patches within question-relevant segments, improving spatial disambiguation. While spatial-temporal modeling is effective, some models without it still perform competitively—most notably CAT (86.10%), which leverages large-scale pretrained vision encoders (ImageBind) and LLaMA2 to infer structure implicitly. However, such models may rely heavily on correlation learned from pretraining, rather than explicit reasoning about visual dynamics. Spatial-temporal models, by contrast, explicitly model the temporal unfolding of gestures and the spatial focus of performer activity—important properties in musical scenes where instrument positions are static but their activation varies over time. These architectural patterns help stabilize attention and reduce confusion when multiple instruments are visually present but only some are active, contributing to more consistent visual QA performance across counting and localization tasks.
@@
Spatial-temporal design is critical for audio-visual QA, where accurate reasoning requires precise temporal and spatial alignment between modalities. Among all Music AVQA categories, audio-visual questions impose the strongest demand on cross-modal synchronization, requiring the model to associate specific acoustic events with their visual sources over time. As shown in Figures 2(f)2(i) and Table 2, models with spatial-temporal components consistently achieve higher accuracy across AV-Existential, AV-Counting, AV-Location, AV-Comparative, and AV-Temporal types. Amuse reaches 82.43% on overall AV questions by leveraging segment-level alignment between synchronized beat-level audio and video inputs and applying cross-modal adapters at each step. PSTP-Net adopts a progressive three-stage pipeline: temporal segment selection, spatial region refinement, and audio-guided attention, resulting in 72.57% AV average. Meerkat further enhances local alignment by explicitly modeling cross-modal transport between audio patches and visual regions, and enforces bounding box constraints for grounding, yielding strong performance on AV-Comparative and AV-Location. In contrast, models without spatial-temporal design—such as MCAN (57.80%), GPT-4o (50.08%), and Qwen2.5-VL (47.75%)—struggle to resolve fine-grained multimodal relationships. While CAT achieves 83.20% AV average through large-scale pretrained encoders, its performance drops on AV-Temporal and AV-Location tasks that require precise temporal ordering or spatial binding. These results support that spatial-temporal designs—especially those involving temporally segmented reasoning, audio-guided spatial focus, and per-frame fusion—enable the model to track which instrument is sounding, when, and in which location, which is critical for answering questions such as “Did the cello on the left play after the drum on the right?”. Without such structure, models tend to conflate co-occurring signals or miss temporally offset actions, leading to lower accuracy in complex cross-modal scenarios.
@@ 7 A Call on Specialized Musical Designs for Music AVQA
Structured musical features (tempo, key, downbeats, chord progressions) provide a coherent framework for aligning audio-visual inputs across time. Symbolic or MIR-derived signals yield interpretable, temporally smooth trajectories that reflect music’s hierarchical organization (phrases, sections, transitions). They abstract from low-level waveform fluctuations and provide a musically meaningful scaffold that persists across genres, tempos, and instrumentation. Integrating them as auxiliary inputs or attention-guiding signals can improve capture of long-range dependencies, maintenance of rhythmic continuity, and resolution of ambiguous instrument interactions, especially in polyphonic or ensemble contexts. This structured conditioning serves as a musical inductive bias, particularly helpful in complex multimodal scenes where overlapping sources challenge bottom-up fusion strategies and where salient events may not be visually or acoustically distinct without temporal alignment cues.
@@
Many Music AVQA questions hinge on implicit causal or temporal relations (e.g., who initiated a phrase, whether an entrance shifted ensemble balance). They often lack step-level supervision, so labels alone do not reveal reasoning paths. We introduce latent reasoning trajectories: structured internal variables representing evolving hypotheses about the musical scene. Rather than mapping inputs to answers, the model infers intermediate latent states (e.g., current lead instrument, rhythmic-intensity change, impending entrances) and updates them as multimodal evidence arrives. Architecturally, this can be realized with hierarchical latent-variable models or recurrent variational modules, with states encoding musical intentions, transitions, and causal flow. These hidden trajectories simulate plausible event sequences, enabling extrapolation and bridging missing links between observed signals. This latent reasoning improves generalization by embedding inductive structure aligned with human inference of musical cause and progression, beyond surface audio-visual co-occurrence.
@@
Some musical questions, especially with temporal or causal dependencies, require sequential sub-decisions to reach the correct answer. For example, “Which instrument enters after the piano stops?” involves (1) detecting the piano stop, (2) detecting subsequent onsets, and (3) selecting the earliest new instrument. Rather than black-box classification, models can be trained to emit intermediate steps via supervised rationales or pseudo-labels from MIR event detection. This approach, akin to chain-of-thought (CoT) in LLMs, improves transparency, supports modular subgoals, and maintains cross-modal alignment. Stepwise supervision highlights failure modes in temporal or semantic inference, providing clearer diagnostics. In music context, CoT chains can include domain-specific steps such as beat alignment, timbre matching, or onset attribution. These interpretable traces support higher accuracy for multi-stage queries and expose reasoning shortcuts and dataset biases.
@@ 8 Conclusion
This paper underscores that successful Music AVQA necessitates specialized designs tailored to the unique demands of musical content: fine-grained audio-visual processing, precise spatial-temporal modeling, and integrated domain knowledge. We call on the research community to: (i) develop more nuanced music understanding benchmarks, and (ii) explore hybrid architectures combining MLLM strengths with music-specific components. The future of effective multimodal AI lies not in a universal approach, but in the thoughtful integration of general capabilities with deep, domain-specific expertise, benefiting music understanding and other complex fields.
Limitations
This paper has several limitations. First, as Music AVQA continues to evolve rapidly, some very recent works may not be fully reflected in this manuscript. Second, although we aim to provide a clear and structured synthesis of existing datasets, methods, and future directions, our organization of the field reflects one possible perspective, and alternative categorizations may also be reasonable. Third, our discussion mainly emphasizes model designs and reasoning patterns, and therefore does not cover aspects such as data collection and human evaluation in equal depth.
Ethical Considerations
All experiments presented in this study were conducted using publicly available datasets and models licensed for academic research purposes. To the best of our knowledge, this work does not present any ethical concerns.
@@ Appendix A Quantitative Comparison on Music AVQA Datasets
We present comprehensive quantitative comparisons of recent state-of-the-art methods on multiple Music AVQA datasets Li et al. (2022); Liu et al. (2024b); Ma et al. (2024), shown in Table 23, and 4. We evaluate the models across a diverse set of question categories, spanning Audio-related, Visual-related, and Audio&Visual-related reasoning tasks. For each dataset, we report accuracy metrics for subcategories such as Counting, Comparative, Location, Existential, and Temporal reasoning, along with average accuracy within each modality and the overall performance.
@@
Amuse Diao et al. (2024) 84.61 82.45 83.58 87.41 84.39 85.84 86.95 85.49 73.01 82.98 83.06 82.43 83.52
Audio Flamingo Kong et al. (2024) - - - - - - - - - - - - -
AVMoE Cheng et al. (2025) - - 77.60 - - 82.70 - - - - - 71.90 75.70
AVSD Schwartz et al. (2019) 72.41 61.90 68.52 67.39 74.19 70.83 81.61 58.79 63.89 61.52 61.41 65.49 67.44
AVSiam Lin and Bertasius (2024) - - - - - - - - - - - - -
AVST Li et al. (2022) 77.78 67.17 73.87 73.52 75.27 74.40 82.49 69.88 64.24 64.67 65.82 69.53 71.59
CAT Ye et al. (2024) - - 84.90 - - 86.10 - - - - - 83.20 84.30
ChatBridge Zhao et al. (2023) - - 28.90 - - 33.10 - - - - - 43.00 78.90
CIGN Mo et al. (2023) - - - - - - - - - - - - -
COCA Lao et al. (2023) 79.94 67.68 75.42 75.10 75.43 75.23 83.50 66.63 69.72 64.12 65.57 69.96 72.33
CONVLSTM Fayek and Johnson (2020) 68.88 63.06 66.73 64.89 58.55 61.68 82.81 55.99 61.30 53.45 54.73 61.75 62.61
CrossMAE Guo et al. (2024) - - - - - - - - - - - - -
DCL Lv et al. (2023) - - - - - - - - - - - - -
DG-SCT Duan et al. (2023) 83.27 64.56 76.34 81.57 82.57 82.08 81.61 72.84 65.91 64.22 67.48 70.56 74.62
EEMC Wang et al. (2024a) - - - - - - - - - - - - -
FCNLSTM Fayek and Johnson (2020) 69.96 61.06 66.67 63.89 58.14 60.98 83.42 56.31 60.28 50.85 56.92 61.46 62.25
GPT-4o Hurst et al. (2024) 65.42 36.07 50.75 72.36 62.30 67.33 56.12 54.84 59.23 37.84 42.35 50.08 54.06
GRU Antol et al. (2015) 71.82 58.90 67.04 66.06 71.82 68.97 81.41 60.30 62.32 56.23 61.89 64.26 66.00
HCRN Le et al. (2020) 70.21 45.62 61.14 62.41 51.51 56.90 52.94 42.07 54.70 50.59 33.33 48.41 52.54
LAST-Att Liu et al. (2024b) 85.71 63.10 - 83.86 83.09 - 76.47 76.20 68.91 65.60 66.75 - 75.45
LAVisH Lin et al. (2023) 75.59 84.13 76.86 77.45 72.91 76.29 71.91 77.52 75.81 76.75 77.62 76.31 76.10
LAViT Yun et al. (2021) 74.36 64.56 70.73 69.39 75.65 72.56 81.21 59.33 64.91 64.22 63.23 66.64 68.93
LSTTA Liu et al. (2023) 81.75 82.04 81.90 81.82 82.23 82.03 83.46 79.11 78.23 78.02 79.32 79.63 81.19
MAVEN Ma et al. (2025) 79.44 54.10 72.79 80.49 93.50 86.99 87.00 66.67 73.85 54.95 68.24 69.94 74.60
MCAN Yu et al. (2019) 75.05 54.58 67.47 68.06 72.15 70.13 81.91 54.15 53.45 52.11 47.21 57.80 62.77
MCCD Ma et al. (2024) 83.87 71.04 79.14 79.78 76.73 78.24 80.87 51.63 71.46 64.67 64.60 67.13 72.20
Meerkat Chowdhury et al. (2025) - - - - - - - 85.70 - 75.98 - - -
OneLLM Han et al. (2024a) - - - - - - - - - - - - 47.60
OPM Wei et al. (2024) - - - - - - - - - - - - 70.80
PSAC Li et al. (2019) 71.33 56.07 65.68 65.89 72.07 69.02 78.59 54.80 63.11 55.96 61.17 62.75 64.92
PSTP-Net Li et al. (2023) 73.97 65.59 70.90 77.15 77.36 77.26 76.18 73.23 71.80 71.19 69.00 72.57 73.52
QaP Liang et al. (2024) - - - - - - - - - - - - -
Qwen2.5-VL Bai et al. (2025) 48.60 55.00 51.80 55.28 53.66 54.47 44.00 52.17 63.57 37.84 41.18 47.75 50.14
RefAtomNet Peng et al. (2024) - - - - - - - - - - - - -
VALOR Liu et al. (2025a) - - 68.70 - - 74.20 - - - - - 75.30 78.90
VAST Chen et al. (2023b) 78.18 67.05 74.06 71.56 76.38 74.00 81.81 64.51 70.80 66.01 63.23 69.54 71.52
VideoLLaMA-2 Cheng et al. (2024) 79.44 52.46 69.64 81.30 82.93 82.11 77.00 63.44 77.69 59.46 64.71 68.98 72.56
VITA Fu et al. (2024) 59.81 45.90 54.76 50.41 34.96 42.68 54.00 49.46 46.92 27.93 41.18 43.74 45.44
@@
Table 2: Comparison with state-of-the-art methods on the Music AVQA Li et al. (2022) test set. We report the accuracy for Audio (Counting, Comparative), Visual (Counting, Location), and Audio-Visual (Existential, Counting, Location, Comparative, Temporal) question types, along with the average accuracy for Audio, Visual, Audio-Visual, and overall. Methods highlighted with a gray background incorporate spatial-temporal designs.
@@
Amuse Diao et al. (2024) 84.76 83.88 84.34 88.15 85.16 86.74 88.30 87.47 78.77 84.41 85.38 85.51 85.16
AVST Li et al. (2022) 81.74 62.11 72.46 79.08 77.64 78.40 72.12 69.03 65.05 63.98 60.57 66.26 71.08
DG-SCT Duan et al. (2023) 83.66 62.47 73.64 82.05 82.97 82.48 83.43 72.70 64.65 64.78 67.34 70.38 74.08
LAST-Att Liu et al. (2024b) 86.03 62.52 - 84.12 84.01 - 76.21 75.23 68.91 65.60 60.60 - 75.44
LAVisH Lin et al. (2023) 84.36 58.57 72.17 83.25 81.46 82.40 73.26 73.45 65.64 64.26 60.82 67.75 72.34
Table 3: Comparison with state-of-the-art methods on the Music AVQA v2.0 Liu et al. (2024b) test set. We report the accuracy for Audio (Counting, Comparative), Visual (Counting, Location), and Audio-Visual (Existential, Counting, Location, Comparative, Temporal) question types, along with the average accuracy for Audio, Visual, Audio-Visual, and overall. Methods highlighted with a gray background incorporate spatial-temporal designs.